# Stability of flocking in the reciprocal two-species Vicsek model: Effects of relative population, motility, and noise

### v1

@@ I Introduction
Natural environments are inherently heterogeneous, where multi-species swarms resemble moving ecosystems Ben-Jacob et al. (2016), and this heterogeneity in mixed populations influences dynamics across scales. Heterogeneity is a natural feature of collective behavior and exists even within a single species due to individual behavioral differences. Examples range from individual fish adjusting their behavior in groups Herbert-Read et al. (2013) to the effect of cell aspect ratio on swarming bacteria Ilkanaiv et al. (2017); Peled et al. (2021). The influence of such individual-level heterogeneity on collective behavior gets further amplified in multi-species systems, where inter-species differences introduce additional complexity. For instance, the ratio of two swarming bacterial species populations has been found to influence dynamics at all scales, from the microscopic speed distribution to mesoscopic vortex sizes and macroscopic colony structure Jose et al. (2022). In mixed-species bacterial swarms, the population ratio can also dictate local segregation Natan et al. (2022). Further examples of heterogeneous systems of self-propelled agents include agents with varying motility Zuo and Wu (2020); Kolb and Klotsa (2020); Pattanayak et al. (2020); Forget et al. (2022); Maity and Morin (2023), diffusivity Pigolotti and Benzi (2014), responsiveness to external cues Book et al. (2017), interparticle interactions Khodygo et al. (2019), temporal characteristics of the heterogeneity Khelfa et al. (2022), and sensitivities to external noise Ariel et al. (2015); Netzer et al. (2019); Ilker and Joanny (2020). Environmental heterogeneity also plays a pivotal role in shaping collective motion across diverse systems. Examples range from bacteria adapting to light cues Wilde and Mullineaux (2017); Frangipane et al. (2018) and active Brownian particles in spatially varying activity landscapes Söker et al. (2021); Auschra et al. (2021); Wysocki et al. (2022) to binary chiral particles under complex environmental noise Huang and Shao (2024), run-and-tumble disks driven through a random obstacle array Reichhardt and Reichhardt (2018) and self-trapping of active particles in disordered media Saavedra and Peruani (2024). Remarkably, topological flocking models maintain long-range order even in spatially heterogeneous environments Rahmani et al. (2021), in contrast to their metric counterparts. These insights highlight the profound role of heterogeneity in governing mesoscale dynamics across natural and synthetic active matter systems.
@@ II Model
?figurename? 1: (color online) Schematic of the different heterogeneities applied on the TSVM. Particles of species A (B) are represented by red (blue) balls. (a) Population heterogeneity: $N_{\rm A}\neq N_{\rm B}$ with $v_{\rm A}=v_{\rm B}$ and $\eta_{\rm A}=\eta_{\rm B}$; (b) Motility heterogeneity: $v_{\rm A}\neq v_{\rm B}$ with $N_{\rm A}=N_{\rm B}$ and $\eta_{\rm A}=\eta_{\rm B}$; (c) Spatial heterogeneity: $v_{\rm A}>v_{\rm B}$ in left region and $v_{\rm A}<v_{\rm B}$ in right region, where the dotted line represents the inter-region interface, $N_{\rm A}=N_{\rm B}$ and $\eta_{\rm A}=\eta_{\rm B}$; (d) Noise heterogeneity: $\eta_{\rm A}\neq\eta_{\rm B}$ with $N_{\rm A}=N_{\rm B}$ and $v_{\rm A}=v_{\rm B}$, i.e. the red particles experience a higher noise amplitude (a “hotter” environment) compared to the blue particles (a “colder” environment).
@@
$\bm{\bar{\sigma}}_{i}^{t}\ =\left(\cos\bar{\theta}_{i}^{t},\sin\bar{\theta}_{i}^{t}\right)\ =\sum_{j\in\mathcal{N}_{i}}J_{ij}\bm{\sigma}_{j}^{t}\ ,$ (1)
@@
$\theta_{i}^{t+\Delta t}\ =\ \bar{\theta}_{i}^{t}\ +\ \eta_{i}\xi_{i}^{t}\,,$ (2)
[…]
@@
[…]
@@ III Simulation details
The TSVM Chatterjee et al. (2023) typically exhibits three phases, a low-density, high-noise gas phase, a low-noise, high-density liquid phase, and an intermediate liquid-gas coexistence region which can further be classified into two categories: (i) PF or ?parallel flocking? state where bands of two species move in the same direction and (ii) APF or ?anti-parallel flocking? state where A and B bands move in the opposite direction. To characterize the collective motion of the A and B species, the following order parameters are introduced Chatterjee et al. (2023):
@@
Let $v_{\pm}=|\bm{v}_{\pm}^{t}|$, then $\langle v_{\pm}\rangle$ are the flocking order parameters, where ()delimited-⟨$\langle$(...$\rangle$ ( … ) ⟩ denotes the steady state time average and the ensemble average over independent runs. The PF and APF states are distinguished by:
@@ IV Results
C. Spatial heterogeneity where particle velocities are space-dependent. In one region of the simulation box, A moves faster than B ($v_{\rm A}>v_{\rm B}$) but in the other region, B has the greater velocity ($v_{\rm A}<v_{\rm B}$). Here also $N_{\rm A}=N_{\rm B}$ and $\eta_{\rm A}=\eta_{\rm B}=\eta$.
@@ IV.1 Population heterogeneity
To characterize the behavior of the PF and APF states separately with $m_{0}$, in Fig. 4, we measure the time-averaged order parameters $\langle v_{s}\rangle$ and $\langle v_{a}\rangle$ averaging only over the ensemble defined by $v_{a}\leq v_{s}$ or only over the ensemble defined by $v_{a}\geq v_{s}$. Fig. 4(a) shows that $\langle v_{a}\rangle$ increases monotonically with $m_{0}$ in the former, and the system is in a PF state, while it remains nearly constant in the latter, where APF behavior dominates. Conversely, Fig. 4(b) shows the opposite trend for $\langle v_{s}\rangle$. The emerging general picture is that the order parameter (e.g., $\langle v_{s}\rangle$) associated with the less prevalent dynamical state (e.g., PF behavior in the $v_{a}\geq v_{s}$ ensemble) approaches that of the dominant state (e.g., $\langle v_{a}\rangle$ representing APF behavior in the same ensemble) as the population heterogeneity $m_{0}$ increases. This convergence signifies a collapse into a single state near $m_{0}\sim 1$, corresponding to the VM limit. In the pure VM limit, $v_{s}$ and $v_{a}$ are equivalent $(v_{s}\simeq v_{a})$, and the system can be described by a single Vicsek order parameter Vicsek et al. (1995). Note that the PF and APF dynamical states are only meaningful when both species form well-defined high-density liquid bands that move either parallel or anti-parallel to each other. When one species becomes significantly more abundant than the other, the concept breaks down, as the minority species can no longer form bands.
@@
?figurename? 6: (color online) Phase diagrams for population heterogeneity. (a) $\eta-m_{0}$ phase diagram for $\rho=2$; (b) $\rho-m_{0}$ phase diagram for $\eta=0.3$. For both cases, the velocity moduli is fixed ($v_{0}=0.5$). Alongside the homogeneous TSVM phases, a single-species flocking state emerges. The boundary lines are included as visual guides.
@@
In summary, extreme population heterogeneity ultimately destroys the PF and APF states into a single unified flocking state (SSF state), at high noise (or low density), where the minority species remains in a gas phase. Although, at low noise (or high density), the minority species still has a directed motion to form an APF liquid state.
@@ IV.2 Motility heterogeneity, or unfriendly ?fast? and ?slow? particles
?figurename? 9: (color online) Order parameters and PF state probability for motility heterogeneity. (a) $\langle v_{a}\rangle$ and (b) $\langle v_{s}\rangle$ in the restricted APF (blue square) and PF (red circle) ensembles versus $\Delta v$ for $\rho=0.5$ and $\eta=0.24$. (c–d) Probability of the PF state ($p_{\rm PF}$) vs $\Delta v$ for (c) varying noise strength $\eta$ [$=0.30$ (circle), $0.35$ (square) and $0.40$ (diamond)] keeping $\rho=2$ fixed and (d) varying particle density $\rho$ [$=3$ (circle), $4$ (square) and $6$ (diamond)] keeping $\eta=0.4$ fixed. Parameters: $v_{0}=0.5$, $L_{x}=256$, and $L_{y}=32$.
@@
In Fig. 10, we present the $\eta-\Delta v$ and $\rho-\Delta v$ phase diagrams for motility heterogeneity, confirming the dominance of the APF state away from $\Delta v=0$. The system remains in a gaseous state at very high noise and low density for all $\Delta v$. As noise decreases or density increases, the system transitions into a liquid-gas coexistence regime, showing PF+APF coexistence for intermediate noise and density values around $\Delta v=0$. Moving further from $\Delta v=0$ along with reducing noise or increasing density, the system first exhibits an APF coexistence state and then eventually enters the APF liquid state at very low noise or very high density.
@@ IV.3 Spatial heterogeneity, or ?activity landscape?
In Fig. 12, we demonstrate how $v_{\rm R}$ and $\zeta$ impact the behavior of the system. We first discuss the impact of $v_{\rm R}$ when the fast and slow regions have an equal size [$\zeta=0.5$, Fig. 12(a–c)]. The system exhibits a PF state when $v_{\rm R}$ is small [$v_{\rm R}=0.1$, Fig. 12(a)] but transitions to an APF state when $v_{\rm R}$ increases [$v_{\rm R}=0.25$, Fig. 12(b)], since the enhanced velocity difference promotes APF behavior (see Sec. IV.2). As A-particles traverse the region I much faster than region II, it leads to wider, more diffuse bands in the fast region and more condensed bands in the slow region. For sufficiently large $v_{\rm R}$ [$v_{\rm R}=0.4$, Fig. 12(c)], A-particles move rapidly through region I, which limits their interaction time with B-particles, and upon entering region II, they slow down significantly and become almost trapped. The average number of A-particles in one region is proportional to the time spent in that region, which is $\zeta L_{x}/(v_{\rm B}+v_{\rm R})$ in region I and $(1-\zeta)L_{x}/(v_{\rm B}-v_{\rm R})$ in region II. The average density then reads
@@
Next, we discuss the case of fast and slow regions with unequal sizes [$\zeta=0.75$, Fig. 12(d–f)]. For $v_{\rm R}=0.1$, the system exhibits an APF state [Fig. 12(d)], in contrast to the corresponding $\zeta=0.5$ case, but recovers the PF state at $v_{\rm R}=0.25$ [Fig. 12(e)]. This suggests that the emergence of parallel flocking in our activity landscape depends on the interplay between $v_{\rm R}$ and $\zeta$, with a larger $\zeta$ requiring a higher $v_{\rm R}$ to sustain the PF state. Similar to Fig. 12(c), for large enough $v_{\rm R}$ [$v_{\rm R}=0.4$, Fig. 12(f)], we observe the trapping of A-particles and the emergence of a vertical APF liquid state in the slow region. However, the trapping is more pronounced for $\zeta=0.75$, due to a narrower slow region, stabilizing even more the vertical APF state since the average density in region II is increased ($\rho_{\rm A}^{\rm II}=\rho_{\rm A}/(1-\zeta)$ when $v_{\rm R}\to v_{\rm B}$).
@@
As we increase $\zeta$, keeping $v_{\rm R}=0.2$ fixed [Fig. 13(f–h)], we observe a non-monotonic behavior of the system concerning the APF state. While an increase in region I $(v_{\rm A}>v_{\rm B})$ encourages parallel flocking, the velocity difference $v_{\rm R}$ also plays a key role. In region I, due to comparatively higher particle velocities $(v_{\rm A}=0.7,v_{\rm B}=0.5)$, more horizontal space $(\zeta L_{x})$ is required for A-particles to catch up, and eventually overtake B-particles. This can lead to A- and B-particles gaining spatial segregation in a PF state on increasing $\zeta$ up to a limit depending upon $v_{\rm R}$, exhibiting PF behavior [Fig. 13(g,h)]. Such overtaking maneuvers can not occur for the simple motility heterogeneity (Sec. IV.2) as one species is consistently faster than the other. Note that, with increasing $\zeta$, the segregation in region II is also decreasing concurrently, discouraging overtakes within the reducing horizontal space $[(1-\zeta)L_{x}]$ and the dominance of the APF behavior is gradually regained [Fig. 13(i, j)].

### v2

@@ I Introduction
Natural environments are inherently heterogeneous, where multi-species swarms resemble moving ecosystems Ben-Jacob et al. (2016), and this heterogeneity in mixed populations influences dynamics across scales. Heterogeneity is a natural feature of collective behavior and exists even within a single species due to individual behavioral differences. Examples range from individual fish adjusting their behavior in groups Herbert-Read et al. (2013) to the effect of cell aspect ratio on swarming bacteria Ilkanaiv et al. (2017); Peled et al. (2021). The influence of such individual-level heterogeneity on collective behavior gets further amplified in multi-species systems, where inter-species differences introduce additional complexity. For instance, the ratio of two swarming bacterial species populations has been found to influence dynamics at all scales, from the microscopic speed distribution to mesoscopic vortex sizes and macroscopic colony structure Jose et al. (2022). In mixed-species bacterial swarms, the population ratio can also dictate local segregation Natan et al. (2022). Further examples of heterogeneous systems of self-propelled agents include agents with varying motility Zuo and Wu (2020); Kolb and Klotsa (2020); Pattanayak et al. (2020); Forget et al. (2022); Maity and Morin (2023), diffusivity Pigolotti and Benzi (2014), responsiveness to external cues Book et al. (2017), interparticle and inter-species interactions Khodygo et al. (2019); Lardet et al. (2025), temporal characteristics of the heterogeneity Khelfa et al. (2022), and sensitivities to external noise Ariel et al. (2015); Netzer et al. (2019); Ilker and Joanny (2020). Environmental heterogeneity also plays a pivotal role in shaping collective motion across diverse systems. Examples range from bacteria adapting to light cues Wilde and Mullineaux (2017); Frangipane et al. (2018) and active Brownian particles in spatially varying activity landscapes Söker et al. (2021); Auschra et al. (2021); Wysocki et al. (2022) to binary chiral particles under complex environmental noise Huang and Shao (2024), run-and-tumble disks driven through a random obstacle array Reichhardt and Reichhardt (2018) and self-trapping of active particles in disordered media Saavedra and Peruani (2024). Remarkably, topological flocking models maintain long-range order even in spatially heterogeneous environments Rahmani et al. (2021), in contrast to their metric counterparts. These insights highlight the profound role of heterogeneity in governing mesoscale dynamics across natural and synthetic active matter systems.
@@ II Model
?figurename? 1: (color online) Schematic of the different heterogeneities applied on the TSVM. Particles of species A (B) are represented by red (blue) balls. (a) Population heterogeneity: $N_{\rm A}\neq N_{\rm B}$ with $v_{\rm A}=v_{\rm B}$ and $\eta_{\rm A}=\eta_{\rm B}$; (b) Motility heterogeneity: $v_{\rm A}\neq v_{\rm B}$ with $N_{\rm A}=N_{\rm B}$ and $\eta_{\rm A}=\eta_{\rm B}$; (c) Spatial heterogeneity: $v_{\rm A}>v_{\rm B}$ in left region and $v_{\rm A}<v_{\rm B}$ in right region, where the dotted line represents the inter-region interface, $N_{\rm A}=N_{\rm B}$ and $\eta_{\rm A}=\eta_{\rm B}$; (d) Noise heterogeneity: $\eta_{\rm A}\neq\eta_{\rm B}$ with $N_{\rm A}=N_{\rm B}$ and $v_{\rm A}=v_{\rm B}$, i.e. the red particles experience a higher noise amplitude (a “hotter” environment) compared to the blue particles (a “colder” environment). In all subsequent snapshots, $L_{x}$ and $L_{y}$ represent horizontal and vertical system sizes, respectively.
@@
$\bm{\bar{\sigma}}_{i}^{t}\;=\;\frac{\displaystyle\sum_{j\in N_{i}}J_{ij}\,\sigma_{j}^{t}}{\bigl{\lvert}\sum_{j\in N_{i}}J_{ij}\,\sigma_{j}^{t}\bigr{\rvert}}\,$ (1)
@@
$\theta_{i}^{t+\Delta t}=\arg\bigl{(}\bm{\bar{\sigma}}_{i}^{t}\bigr{)}+\eta_{i}\,\xi_{i}^{t}\ ,$ (2)
where $\xi_{i}^{t}$ is a scalar noise uniformly distributed in $\pi,\pi]$ and uncorrelated for all sites and times: $\langle\xi_{i}^{t}\rangle\ =\ 0$, and $\langle\xi_{i}^{t}\xi_{j}^{s}\rangle\sim\delta_{ts}\delta_{ij}$, and $\eta_{i}$ is the parameter controlling the noise strength. $\eta_{i}=\eta_{\rm A}$ for A particles and $\eta_{i}=\eta_{\rm B}$ for B particles.
@@
The model parameters include the species densities $\rho_{s}=N_{s}/L_{x}L_{y}$ (for $s\in{\rm\{A,B\}}$), the noise strengths $\eta_{s}$, and the velocity moduli $v_{s}$. For simplicity, we consider the following when these parameters are uniform across species: if all particles share the same speed, we set $v_{\rm A}=v_{\rm B}=v_{0}$, if the noise strengths are identical, we take $\eta_{\rm A}=\eta_{\rm B}=\eta$, and if the species densities are equal, we define $\rho_{\rm A}=\rho_{\rm B}=\rho/2$, where $\rho=N/L_{x}L_{y}$ is the total particle number density of the system. We mostly consider a rectangular simulation box of high aspect ratio $L_{x}/L_{y}\ =\ 8$, $R=1$, and $\Delta t=1$, unless stated otherwise.
@@ III Simulation details
The TSVM Chatterjee et al. (2023) typically exhibits three phases: a low-density, high-noise gas phase, a low-noise, high-density liquid phase, and an intermediate liquid-gas coexistence region which can further be classified into two categories: (i) PF or ?parallel flocking? state where bands of two species move in the same direction and (ii) APF or ?anti-parallel flocking? state where A and B bands move in the opposite direction. To characterize the collective motion of the A and B species, the following order parameters are introduced Chatterjee et al. (2023):
@@
[…]
@@ IV Results
C. Spatial heterogeneity where particle velocities are space-dependent. In one region of the simulation box, A moves faster than B ($v_{\rm A}>v_{\rm B}$), but in the other region, B has the greater velocity ($v_{\rm A}<v_{\rm B}$). Here also $N_{\rm A}=N_{\rm B}$ and $\eta_{\rm A}=\eta_{\rm B}=\eta$.
@@ IV.1 Population heterogeneity
It is important to highlight that the microphase-separated band configurations displayed in Fig. 2 correspond to stable steady states that persist over long time scales. For each value of $m_{0}$, once the system undergoes phase separation, both the number and the form of the bands remain essentially constant, with no observable change up to times of at least $t=10^{7}$ (see Appendix A). This indicates that the system undergoes minimal, if any, coarsening after the initial formation of bands. Although the resulting structures may resemble smectic order due to the apparent regularity in band spacing, significant fluctuations in inter-band distances preclude any sustained long-range translational order. Similar to the standard Vicsek model (VM), the TSVM displays giant number fluctuations (GNF) Chatterjee et al. (2023), which are fundamentally incompatible with smectic order Chen and Toner (2013); Adhyapak et al. (2013). These fluctuations disrupt crystalline arrangements and inhibit translational symmetry breaking. Consequently, even if quasi-periodic band spacing may appear over certain time intervals or regions, the system retains a dynamic, fluctuating banded structure rather than forming a true smectic or crystalline phase (see Appendix A).
@@
To characterize the behavior of the PF and APF states separately with $m_{0}$, in Fig. 4, we measure the time-averaged order parameters $\langle v_{s}\rangle$ and $\langle v_{a}\rangle$, averaging only over the ensemble defined by $v_{a}\leq v_{s}$ or only over the ensemble defined by $v_{a}\geq v_{s}$. Fig. 4(a) shows that $\langle v_{a}\rangle$ increases monotonically with $m_{0}$ in the former, and the system is in a PF state, while it remains nearly constant in the latter, where APF behavior dominates. Conversely, Fig. 4(b) shows the opposite trend for $\langle v_{s}\rangle$. The emerging general picture is that the order parameter (e.g., $\langle v_{s}\rangle$) associated with the less prevalent dynamical state (e.g., PF behavior in the $v_{a}\geq v_{s}$ ensemble) approaches that of the dominant state (e.g., $\langle v_{a}\rangle$ representing APF behavior in the same ensemble) as the population heterogeneity $m_{0}$ increases. This convergence signifies a collapse into a single state near $m_{0}\sim 1$, corresponding to the VM limit. In the pure VM limit, $v_{s}$ and $v_{a}$ are equivalent $(v_{s}\simeq v_{a})$, and the system can be described by a single Vicsek order parameter Vicsek et al. (1995). Note that the PF and APF dynamical states are only meaningful when both species form well-defined high-density liquid bands that move either parallel or anti-parallel to each other. When one species becomes significantly more abundant than the other, the concept breaks down, as the minority species can no longer form bands.
@@
?figurename? 6: (color online) Phase diagrams for population heterogeneity. (a) $\eta-m_{0}$ phase diagram for $\rho=2$; (b) $\rho-m_{0}$ phase diagram for $\eta=0.3$. For both cases, the velocity modulus is fixed ($v_{0}=0.5$). Alongside the homogeneous TSVM phases, a single-species flocking state emerges. The boundary lines are included as visual guides.
@@
In summary, strong population heterogeneity ultimately eliminates the PF and APF states, leading to a single unified flocking state (SSF) at high noise (or low density), where the minority species remains in a disordered gas phase. However, at low noise (or high density), the minority species continues to exhibit directed motion, forming an APF-like liquid state.
@@ IV.2 Motility heterogeneity, or unfriendly ?fast? and ?slow? particles
?figurename? 9: (color online) Order parameters and PF state probability for motility heterogeneity. (a) $\langle v_{a}\rangle$ and (b) $\langle v_{s}\rangle$ in the restricted APF (blue square) and PF (red circle) ensembles versus $\Delta v$ for $\rho=0.5$ and $\eta=0.24$. (c–d) Probability of the PF state ($p_{\rm PF}$) vs $\Delta v$ for (c) varying noise strength $\eta$ [=0.30=0.30 (circle), $0.35$ (square) and $0.40$ (diamond)] keeping $\rho=2$ fixed and (d) varying particle density $\rho$ [=3=3 (circle), 44 (square) and 66 (diamond)] keeping $\eta=0.4$ fixed. Parameters: $v_{0}=0.5$, $L_{x}=256$, and $L_{y}=32$.
@@
In Fig. 10, we present the $\eta-\Delta v$ and $\rho-\Delta v$ phase diagrams for motility heterogeneity, confirming the dominance of the APF state away from $\Delta v=0$. The system remains in a gaseous state at very high noise and low density for all $\Delta v$. As noise decreases or density increases, the system transitions into a liquid-gas coexistence regime, showing PF+APF coexistence for intermediate noise and density values around $\Delta v=0$. Moving further from $\Delta v=0$ along with reducing noise or increasing density, the system first exhibits an APF coexistence state and then eventually enters the APF liquid state at very low noise or very high density. It is worth noting that, although reversing the sign of motility heterogeneity, $\Delta v\to-\Delta v$, simply swaps the roles of the fast and slow species, the phase diagrams in Fig. 10 are not symmetric under this transformation. This asymmetry arises because $\Delta v$ is varied while keeping $v_{\rm B}=0.5$ fixed, so that $v_{\rm A}<0.5$ for negative $\Delta v$ and $v_{\rm A}>0.5$ for positive $\Delta v$. For fixed noise $\eta$ and density $\rho$, the steady state for $\Delta v<0$ is generally less ordered than that for $\Delta v>0$.
@@ IV.3 Spatial heterogeneity, or ?activity landscape?
In Fig. 12, we demonstrate how $v_{\rm R}$ and $\zeta$ impact the behavior of the system. We first discuss the impact of $v_{\rm R}$ when the fast and slow regions have an equal size $\zeta=0.5$, Fig. 12(a–c)]. The system exhibits a PF state when $v_{\rm R}$ is small $v_{\rm R}=0.1$, Fig. 12(a)] but transitions to an APF state when $v_{\rm R}$ increases $v_{\rm R}=0.25$, Fig. 12(b)], since the enhanced velocity difference promotes APF behavior (see Sec. IV.2). As A-particles traverse the region I much faster than region II, it leads to wider, more diffuse bands in the fast region and more condensed bands in the slow region. For sufficiently large $v_{\rm R}$ $v_{\rm R}=0.4$, Fig. 12(c)], A-particles move rapidly through region I, which limits their interaction time with B-particles, and upon entering region II, they slow down significantly and become almost trapped. The average number of A-particles in one region is proportional to the time spent in that region, which is $\zeta L_{x}/(v_{\rm B}+v_{\rm R})$ in region I and $(1-\zeta)L_{x}/(v_{\rm B}-v_{\rm R})$ in region II. The average density then reads
@@
Next, we discuss the case of fast and slow regions with unequal sizes $\zeta=0.75$, Fig. 12(d–f)]. For $v_{\rm R}=0.1$, the system exhibits an APF state [Fig. 12(d)], in contrast to the corresponding $\zeta=0.5$ case, but recovers the PF state at $v_{\rm R}=0.25$ [Fig. 12(e)]. This suggests that the emergence of parallel flocking in our activity landscape depends on the interplay between $v_{\rm R}$ and $\zeta$, with a larger $\zeta$ requiring a higher $v_{\rm R}$ to sustain the PF state. Similar to Fig. 12(c), for large enough $v_{\rm R}$ $v_{\rm R}=0.4$, Fig. 12(f)], we observe the trapping of A-particles and the emergence of a vertical APF liquid state in the slow region. However, the trapping is more pronounced for $\zeta=0.75$, due to a narrower slow region, stabilizing even more the vertical APF state since the average density in region II is increased ($\rho_{\rm A}^{\rm II}=\rho_{\rm A}/(1-\zeta)$ when $v_{\rm R}\to v_{\rm B}$).
@@
As we increase $\zeta$, keeping $v_{\rm R}=0.2$ fixed [Fig. 13(f–h)], we observe a non-monotonic behavior of the system concerning the APF state. While an increase in region I $(v_{\rm A}>v_{\rm B})$ encourages parallel flocking, the velocity difference $v_{\rm R}$ also plays a key role. In region I, due to comparatively higher particle velocities $(v_{\rm A}=0.7,v_{\rm B}=0.5)$, more horizontal space $(\zeta L_{x})$ is required for A-particles to catch up, and eventually overtake B-particles. This can lead to A- and B-particles gaining spatial segregation in a PF state on increasing $\zeta$ up to a limit depending upon $v_{\rm R}$, exhibiting PF behavior [Fig. 13(g,h)]. Such overtaking maneuvers can not occur for the simple motility heterogeneity (Sec. IV.2) as one species is consistently faster than the other. Note that, with increasing $\zeta$, the segregation in region II is also decreasing concurrently, discouraging overtakes within the reducing horizontal space $(1-\zeta)L_{x}]$ and the dominance of the APF behavior is gradually regained [Fig. 13(i, j)].